OpenStudy (anonymous):

find the perimeter o fthe square the lenght is 8x-10 the width is 6x

OpenStudy (anonymous):

square means sides are the same length

OpenStudy (anonymous):

so 8x-10=6x

OpenStudy (anonymous):

so how do u work it out

OpenStudy (anonymous):

2(4x-5)=6x

OpenStudy (anonymous):

4x-5=3x

OpenStudy (anonymous):

-5=-x

OpenStudy (anonymous):

5=x

OpenStudy (anonymous):

Now plug x back in...

OpenStudy (anonymous):

6x=6(5)=30

OpenStudy (anonymous):

perimeter is the sum of all the sides, so do 30+30+30+30 (or 30*4) to get 1200

OpenStudy (anonymous):

so whats x 5

OpenStudy (anonymous):

x5?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

x is five right

OpenStudy (anonymous):

yes, x= 5

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

welcome